How do I solve 4(x+3)^2-2=62?
Answer by jim_thompson5910(35256)


4(x+3)^2-2=62

4(x^2+6x+9)-2=62

4x^2+24x+36-2=62

4x^2+24x+36-2-62=0

4x^2+24x-28=0

4(x^2+6x-7)=0

x^2+6x-7=0/4

x^2+6x-7=0

(x-1)(x+7)=0

x-1 = 0 or x+7 = 0

x = 1 or x = -7

So the solutions are x = 1 or x = -7

If you need more help, feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim